40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 25 17:42:11 2021, Anchal, Paco, Summary, ALS, ALS Single Arm Noise Budget ALS_IR_b.svgALS_Single_Arm_IR.pdf
    Reply  Thu May 27 11:03:15 2021, Anchal, Paco, Summary, ALS, ALS Single Arm Noise Budget ALS_Single_X_Arm_IR.pdfALS_OOL_with_Ref.pdf
       Reply  Fri May 28 17:32:48 2021, Anchal, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat SingleArmActCalwithIRALSBeat.pdfstateSpaceModel.zip
          Reply  Tue Jun 1 16:55:32 2021, Anchal, Paco, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat SingleArmActCalwithIRALSBeat-1306624785.pdf
             Reply  Tue Jun 8 11:40:53 2021, Anchal, Paco, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat ITMX_Cal_Noise_Spectrum_1307143423.pdf
                Reply  Wed Jun 9 18:29:13 2021, Anchal, Paco, Summary, ALS, Check for saturation in ITMX SOS Driver 
                Reply  Fri Jul 9 15:39:08 2021, Anchal, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat [Correction] ITMX_calibration_With_ALS_Beat.pdf
Message ID: 16164     Entry time: Thu May 27 11:03:15 2021     In reply to: 16161     Reply to this: 16168
Author: Anchal, Paco 
Type: Summary 
Category: ALS 
Subject: ALS Single Arm Noise Budget 

Here's an updated X ARM ALS noise budget.

Things to remember:

  • Major mistake we were making earlier was that we were missing the step of clicking  'Set Phase UGF' before taking the measurement.
  • Click the clear phase history just before taking measure.
  • Make sure the IR beatnotes are less than 50 MHz (or the left half of HP8591E on water crate). The DFD is designed for this much beatnote frequency (from Gautum).
  • We took this measurement with old IMC settings.
  • We have saved a template file in users/Templates/ALS/ALS_outOfLoop_Ref_DQ.xml . This si same as ALS_outOfLoop_Ref.xml except we changed all channels to _DQ.

Conclusions:

  • Attachment 1 shows the updated noisebudget. The estimated and measured RMS noise are very close to eachother.
  • However, there is significant excess noise between 4 Hz and 200 Hz. We're still thinking on what could be the source of these.
  • From 200 Hz to about 3 kHz, the beatnote noise is dominated by AUX residual frequency noise. This can be verified with page 2 of Attachment 2 where coherence between AUX PDH Error signal and BEATX signal is high.
  • One mystery is how the measured beatnote noise is below the residual green laser noise above 3 kHz. Could this be just because the phase tracker can't measure noise above 3kHz?
  • We have used estimated open loop transfer function for AUX from poles/zeros for uPDH box used (this was done months ago by me when I was working on ALS noise budget from home). We should verify it with a fresh OLTF measurement of AUX PDH loop. That's next on our list.
Attachment 1: ALS_Single_X_Arm_IR.pdf  150 kB  Uploaded Thu May 27 12:09:49 2021  | Hide | Hide all
ALS_Single_X_Arm_IR.pdf
Attachment 2: ALS_OOL_with_Ref.pdf  194 kB  Uploaded Thu May 27 12:12:06 2021  | Hide | Hide all
ALS_OOL_with_Ref.pdf ALS_OOL_with_Ref.pdf ALS_OOL_with_Ref.pdf ALS_OOL_with_Ref.pdf
ELOG V3.1.3-